OpenStudy (anonymous):

This word problem has got me completely stuck: A ship is just offshore of New York City. A sighting is taken of the Statue of Liberty, which is about 305 feet tall. If the angle of elevation to the top of the statue is 22°, how far is the ship from the base of the statue?

OpenStudy (anonymous):

It helps to draw a picture..

OpenStudy (anonymous):

I would advice drawing a diagram. The 305 feet tall refers to the opposite side of a right-angled triangle. Which angle is the angle of elevation? What side of the triangle are you trying to find if you want to know how far the ship is from the base of the statue?

OpenStudy (anonymous):

ive tried 305cos22 because i thought i was looking for the adjacent side of the angle of elevation and my hyp. is 305 and just cross multiply but I keep getting the wrong answer.

OpenStudy (anonymous):

The 305 feet refers to the \(opposite\) side of the the triangle, and not the hypotenuse. How would you go about solving it with this piece of information?

OpenStudy (anonymous):

Here's my terrible rendition.

myininaya (myininaya):

i agree with polpak's drawing

OpenStudy (anonymous):

so tan which is opp/adj and then tan22= 305/x

OpenStudy (anonymous):

\[d=305 \text{ Cot}[22 {}^{\circ}]\]\[d=755 \text{ feet}\]

OpenStudy (anonymous):

oh oh cot! i meant

myininaya (myininaya):

you can do tan

myininaya (myininaya):

solve for x in your equation

OpenStudy (anonymous):

okay, thank you!! (:

OpenStudy (amistre64):

we need a picture that is 1:1 scaled ;)

OpenStudy (anonymous):

\[h(cos\ 22^\circ) = x\]\[h(sin\ 22^\circ) = 305\]\[\implies \frac{x}{305} = \frac{(cos\ 22^\circ) }{(sin\ 22^\circ) }\]\[\implies x =305\frac{(cos\ 22^\circ) }{(sin\ 22^\circ) } \]

myininaya (myininaya):

or x=305cot(22) or x=305/tan(22) all have same meaning

OpenStudy (anonymous):

bah, shoulda used r. h is ambiguous.. I meant it to be hypotenuse but it could be interpreted wrongly as height.

OpenStudy (anonymous):

Yeah, but I hate trying to remember all that other nonsense, cot, etc. easier just to use sines and cosines usually.

myininaya (myininaya):

cot=adjacent/opposite tan= opposite/adjacent

myininaya (myininaya):

sin=opp/hyp cos=adj/hyp tan=sin/cos=opp/adj

OpenStudy (anonymous):

Yeah I can figure it out, but I hate trying to remember all of those relationships.

OpenStudy (anonymous):

Easier for me to remember the basic one, and derive the rest as needed.

myininaya (myininaya):

but i suppose we only really need sin and cos

myininaya (myininaya):

all the trig relations can be written in terms of sine and cosine

OpenStudy (amistre64):

tan(22) = 305/d d = 305/tan(22) is what i get

myininaya (myininaya):

thanks for using derive right amistre i mean polpak

myininaya (myininaya):

its wait i got too :)